A plane takes off from a runway and climbs at at angle of elevation of 15 degrees.
a. After the plane has traveled one mile (5280 feet) how far has the plane traveled horizontally? Round answer to nearest foot.
b.After the plane has traveled one mile (5280 feet) how far has the plane traveled vertically? Round your answer to the nearest foot.

Answer by TimothyLamb(4379)

set up a right triangle with the airplane at the top of the short leg:
---
third angle C:
C = 180 - 90 - 15
C = 75
---
a:
use law of sines:
c/sinC = h/sinH
c = sinC*h/sinH
c = sin(75)*5280/sin(90)
c = 5100.0883628
the airplane traveled 5100 feet horizontally
---
b:
use law of sines:
b/sinB = h/sinH
b = sinB*h/sinH
b = sin(15)*5280/sin(90)
b = 1366.56455814
the airplane traveled 1367 feet vertically
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---